# Evaluation of LLMs for mathematical problem solving

### v1

@@ 6 Conclusion
This study comprehensively evaluated the performance of GPT-4o, DeepSeek-V3, and Gemini-2.0 in mathematical problem-solving tasks in five dimensions, covering three types of datasets, including basic (GSM8K), intermediate (MATH500), and advanced (UNSW). This study presents an automatic scoring framework to evaluate LLMs on mathematical problems featuring multidimensional qualitative analysis. It not only reveals the performance differences between the current mainstream models in mathematical reasoning tasks but also verifies the key role of cue design and structured scoring mechanisms in improving the interpretability of the models.
@@ 7.1 Examples from mathematical problems
(a) MATH5165 Class Test 1
(b) MATH5335 Final Exam 2013
(c) MATH5905 Assignment 1
Figure 6: Selected examples from the UNSW dataset.
7.2 Sample Question and LLM Answers
Figure 7: Sample questions and correct answers.
Figure 8: LLM solutions to mathematical problems.

### v2

@@ 5 Conclusion
This study comprehensively evaluated the performance of GPT-4o, DeepSeek-V3, and Gemini-2.0 in mathematical problem-solving tasks in five dimensions, covering three types of datasets, including basic (GSM8K), intermediate (MATH500), and advanced (University). This study presents an automatic scoring framework to evaluate LLMs on mathematical problems featuring multidimensional qualitative analysis. It not only reveals the performance differences between the current mainstream models in mathematical reasoning tasks but also verifies the key role of cue design and structured scoring mechanisms in improving the interpretability of the models.